Evaluate α given that α= [log(e2elog2)/2]1/2 .


1
Expert's answer
2021-08-09T07:57:12-0400

"\alpha =[log(e^2e^{log2})/2]^{1/2}=[\frac{log(e^2 2)}{2}]^{1/2} = [\frac{log(e^2)+log 2)}{2}]^{1/2}"


"= [\frac{2+log 2)}{2}]^{1/2}=[1+\frac{1}{2}log 2]^{1/2}=\sqrt{1+log(\sqrt 2)}"